Status Report

NASA ISS On-Orbit Status 31 May 2008

By SpaceRef Editor
May 31, 2008
Filed under , , ,
NASA ISS On-Orbit Status 31 May 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – a light-duty but long day for CDR Volkov, FE-1 Kononenko & FE-2 Reisman.

STS-124/Discovery (ISS-1J) lifted off right on time (5:02pm EDT) with all systems performing nominally, for ISS rendezvous on Monday (6/2), to dock at ~1:54pm EDT. At launch, the ISS was off Halifax/Canada, at 42.8 deg N Lat, 57.6 deg W Long. The Orbiter is carrying the seven-member crew of CDR Mark Kelly, PLT Ken Ham, MS1 Karen Nyberg, MS2 Ron Garan, MS3 Mike Fossum, MS4 Akihiko Hoshide & MS5 Greg Chamitoff. Chamitoff will replace ISS Flight Engineer 2 Garrett Reisman who returns on 6/14 (nominal) with STS-124. STS-124 is the 123rd space shuttle flight, the 35th flight for Discovery, the 26th flight to the station and the third Shuttle flight in 2008. Its primary payload, the largest so far, is the 32,000-lbs, 36.7-ft long JPM (Japanese Pressurized Module) with its RMS (Remote Manipulator System). We are off to another great mission! [The eighth crewmember on board is a stow-away: Buzz Lightyear.]

For the ISS crew, who watched the liftoff on their Ku-band/TV uplink, the new wake/sleep cycle shift begins tonight with a late bedtime of 10:00pm EDT (wake-up this morning: 2:00am). [Wake/sleep tomorrow (Sunday): 6:32am-10:02pm; on Monday (docking day): 6:32am-10:02pm.]

Volkov, Kononenko & Reisman performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, the crew also performed regular maintenance inspection & cleaning on fan grilles in the FGB (FS5 & FS6, TsV2), SM (VPkhO & VPrK), DC1 (V3) and in the COL (Columbus Orbital Laboratory).

Garrett Reisman stopped sampling mode on the ANITA (Analyzing Interferometer for Ambient Air) payload, exited its software and powered down the ANITA laptop. Later, at ~2:00pm, ANITA was turned back on.

The FE-2 then supported SAMS (Space Acceleration Measurement System) ground commanding by activating its ICU (Interim Control Unit) in Rack LAP2/Drawer 1. [SAMS acquires acceleration data relevant to characterization of the dynamic vibratory environment on the ISS.]

In preparation for Discovery’s arrival on 6/2, the crew jointly conducted a refresher review of RPM (R-Bar Pitch Maneuver) documentation. [The review prepared the crewmembers for the bottom-side mapping of the Orbiter at its arrival on 6/2. During the RPM at ~600 ft from the station, the “shooters” have only ~90 seconds for taking high-resolution digital photographs of all tile areas and door seals on Discovery, to be downlinked for launch debris assessment. Thus, time available for the shooting will be very limited, requiring great coordination between the two headset-equipped photographers and the Shuttle pilot.]

Performing “get-aheads” for JPM “Kibo” laboratory arrival, FE-2 Reisman worked in the JLP (JEM Logistics Module) after activating the JLP MKAM (Minimum Keep Alive Monitor) fan and temporarily moving the JLP’s PBA (Portable Breathing Apparatus) from its location COL1PF in the COL (Columbus Orbital Laboratory) to the JLP endcone (JLP1F3). [For the upcoming 1J activities, specialists have developed a “PFE/PBA Musical Chairs” plan for moving these fire safety devices around the station’s interior as required for quick access.]

Reisman’s work in the JLP was to retrieve two panels for the JPM from behind payload racks, which he had to rotate down for access and later return to the upright position, a ~10-min task each time. [For a hard decorative panel, stowed at JLP1SA2, Garrett unlocked & rotated the Ryutai rack, for a soft protective panel, at JLP1FP2, the EPS-1 (Electrical Power Systems 1) rack. Ryutai (“fluid”) is a Japanese multipurpose experiment/payload rack system to support the FPEF (Fluid Physics Experiment Facility), SCOF (Solution Crystallization Observation Facility), PCRF (Protein Crystallization Research Facility) and the IPU (Image Processing Unit) by providing structural interfaces, power, data, cooling, water and other items needed to operate science experiments in micro-G.]

The FE-2 then removed a front bracket from the JLP WS (Work Station) Rack and bundled it together with the two panels. Afterwards, he deactivated the JLP MKAM fan and returned the PBA to the COL1PF location.

In the SM, Volkov completed the routine maintenance of the SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. [Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.]

The FE-2 had another hour set aside for more personal departure preparations. [Garrett returns to Earth with STS-124 on 6/14, departing from the ISS on 6/11 (FD12).]

The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), RED resistive exercise device (CDR, FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

Afterwards, Garrett transferred the crew’s exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~10:15am EDT, Kononenko had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

Working off the discretionary “if time permits” task list, Sergey Volkov completed another EKON KPT-3 session, making observations and taking aerial photography of environmental conditions (water contamination in the Kerch Strait) for Russia’s Environmental Safety Agency (EKON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.

Also off the Russian suggestions list, Oleg Kononenko performed a session of the Russian GFI-8 "Uragan" (hurricane) earth-imaging program, using the D2X to take telephotos. [Uplinked target zones were the Pamir mountain range (glacier tongues located to the west of the largest Fedchenko glacier), large glaciers of Tien-Shan located to the east of lake Issyk-Kul, oil slicks in the Caspian Sea, condition of drying out shores of the Aral sea, logging sites and burnt out forest, the Caucasus mountains forest cover condition from the coast to Stavropol plain, the Volga-Akhtubinsk flood plain situation during hydropower plant water spillover (general view and close-up shots).]

WDS Update: An updated Water Delivery System “cue card” was uplinked for the crew’s reference. [The new card (17-0002H) lists 32 CWCs (~1182.2 L total) for the four types of water identified on board: technical water (619.6 L, for Elektron, flushing, hygiene), potable water (534.8 L), condensate water (23 L), waste/EMU dump and other (5.8 L). Of the 32 containers, 11 CWCs with technical water (423.2 L) can only be used for flushing due to Wautersia bacteria contamination, and 2 CWCs with potable water (88.7 L) are also not cleared for use except for flushing.]

ASU Toilet System Update: Detailed instructions were uplinked to the crew on the use of US HRF UCDs (Human Research Facility/Urine Collector Devices) with the Russian Wring Receptacle Assembly using a special adapter and RSQD (Rubber Stopper/Quick Disconnect). After the urine transfer, the Wring Receptacle will be connected to the EDV container for disposal. Wring collectors and UCDs are carried on STS-124, along with a replacement pump for the Russian ASU toilet system’s air/water separator (MNR-NS).

Weekly Science Update (Expedition Seventeen — Week 6)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): In progress.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally, although the platform is rather cold due to unfavourable Beta angles. A DHPU (Data Handling Processing Unit) software patch is currently developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. – DEBIE-2: Link error still under investigation; – DOSTEL: On-going science acquisition. Mode-2 (faster acquisition sampling rate) was used from 5/27 to 5/29; – EuTEMP: Currently inactive as planned; – EVC: Inactive, and could not be powered up due to very low temperature; – EXPOSE: On-going science acquisition; – FIPEX: A science acquisition script has run nominally until 5/26. On 5/28, a new script has been started almost nominally – there is a minor glitch with “fast measure” commands; – MEDET: Since 5/18, the instrument had to be switched off because it reached the low temperature limit. On 5/28, MEDET was powered back on is now acquiring science; – PLEGPAY: Instrument is powered on, but not in science acquisition mode; – TRIBOLAB: The POD#3 (Pin On Disk run#3) was restarted nominally on 5/17 and has been running nominally since then.

FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. Further troubleshooting activities will be required with CEM-U (Upper) Optical Module.

GEOFLOW: Further troubleshooting activities with CEM-U (Upper) Optical Module are required. A new lamp set is manifested on STS-124/1J. GEOFLOW start is pending further FSL troubleshooting, and remaining commissioning activities (check of FSL optical modes). This will not occur prior to 1J flight.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Garrett, thanks for completing your last Actiwatch Download/Initialization session. Your Sleep Actiwatch that you are wearing will be doffed and stowed on the Shuttle with the other Sleep subjects’ Actiwatches. Any additional sleep logging you perform between now and the Shuttle undocking is welcome. Thanks and enjoy the trip home.”

SOLAR (Solar Monitoring Observatory): The last Sun observation period has closed on 5/10. On 5/28, SOLAR was ground commanded to test its CPD (Coarse Pointing Device) and get additional telemetry from the end switches. The collected data will support investigation of the so-called “Zero-Procedure” anomaly. The next observation window will start on 6/3. – SOVIM: Waiting for the Sun; – SOLSPEC: Waiting for the Sun; – SOLACES: Waiting for the Sun.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

3-D SPACE : First session with Inc17 FE-2 is currently planned for 6/13.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): On 5/15, a caution signal was encountered with the BIOLAB Smoke Detector while activating BIOLAB from ground after the Reference ECs insertion to Rotor A crew activity. Engineering teams are analyzing the telemetry and are preparing a recovery plan. On 5/27, the BIOLAB Rotor B Locking Pin actuator was exchanged successfully by Garrett. After that, it was confirmed that the Rotor B (stuck since 1E Stage) can be manually rotated again. Unfortunately, BIOLAB could not be activated from the ground yet, pending resolution of the Smoke Detector issue. After the 1J flight, the remaining WAICO #1 ECs will be disposed of and replaced by six Reference ECs on Rotor B.

CEO (Crew Earth Observations): Through 5/22 the ground has received a total of 1,780 of ISS/CEO images for review and cataloging. “No new imagery in response to our target list request was received this past week. We fully appreciate your demanding schedule in preparation for next week. Your interesting, high-oblique view of the southeastern Hawaiian Islands will be published in NASA/GSFC’s Earth Observatory website this weekend. Your view illustrated the emission and spread of caustic, volcanic fog – known as vog – that occasionally plagues islanders. It also showed cloud formations that result from the impact of these isolated islands on the regional trade winds and local land/sea breezes. Very nice educational picture! We will resume CEO operations and reporting after STS-124.”

CEO (Crew Earth Observation) photo targets uplinked for today were Yellow River Delta, China (weather was predicted to be clear over this dynamic delta, which historically has changed its extent and morphology on almost a daily basis. ISS had a nadir pass over the center of the delta region. High-resolution imagery of the delta region was requested to map its current configuration), and Lake Nasser, Toshka Lakes, Egypt (looking to the left of track for the man-made Toshka Lakes, located to the west of Lake Nasser. Imagery of the current shorelines of the Lakes was requested in order to track changes in water level and extent).

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

ISS Orbit (as of this morning, 9:50am EDT [= epoch]):
Mean altitude — 340.0 km
Apogee height — 343.7 km
Perigee height — 336.4 km
Period — 91.34 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005388
Solar Beta Angle — 40.9 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.76
Mean altitude loss in the last 24 hours — 72 m
Revolutions since FGB/Zarya launch (Nov. 98) — 54590

Significant Events Ahead (all dates Eastern Time, some changes possible.):
06/02/08 – FD3 – STS-124/Discovery/1J docking (1:54pm)

— STS-124 docked timeline:

  • 6/03 – FD4 – EVA-1 (11:30am, 6.5 hrs.), OBSS transfer, JPM prep, S3/S4 SARJ TBA install, JPM install; Fossum/Garan (EV1/EV2)
  • 6/04 – FD5 – JPM Vestibule outfitting, Reisman/Chamitoff handovers
  • 6/05 – FD6 – EVA-2 (11:30am, 7 hrs), JPM outfit (JTVE install, JRMS cvr remv), S1 NTA prep, CP 9 ETVCG retrv
  • 6/06 – FD7 – JLP relocate to JPM; JLP Vestibule leak check; Focused inspection
  • 6/07 – FD8 – JLP Vestibule outfitting; CP9 ETVCG TVCIC R&R
  • 6/08 – FD9 – EVA-3 (10:30am, 7 hrs), S1 NTA R&R, compl JPM outfit (RMS cvr remv), P1 CP9 ETVCG install
  • 6/09 – FD10 – JRMS checkouts, JLP Vestibule outfitting, A/L BCM R&R
  • 6/10 – FD11 – Sayonara (~4:00pm), hatch close (~4:30pm)
  • 6/11 – FD12 – Undocking (~7:33am); Greg remains, Garrett leaves; OBSS survey/inspection
  • 6/12 – FD13 – Mostly off-duty
  • 6/13 – FD14 – Stowing; deorbit preps
  • 6/14 – FD15 – Deorbit burn

06/14/08 — STS-124/Discovery landing (KSC: ~11:02am EDT, nominal)
07/10/08 — Russian EVA-20 (7/10-11)
09/05/08 — ATV1 undocking
09/09/08 — Progress M-64/29P undocking (from DC1)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (at DC1 nadir)
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (SM aft port)
10/23/08 — Soyuz TMA-12/16S undocking (FGB nadir)
11/03/08 — Soyuz TMA-13/17S relocation (from SM aft to FGB nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
12/04/08 — STS-119/Discovery/15A launch – S6 truss segment
12/06/08 — STS-119/Discovery/15A docking
12/15/08 — STS-119/Discovery/15A undocking
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/??/09 — Six-person crew on ISS (following Soyuz 18S-2 docking)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.